OpenStudy (anonymous):

simplafiy 5^-4 ------- 5^1

OpenStudy (unklerhaukus):

\[\frac{5^{-4}}{5^1}\]

OpenStudy (anonymous):

it was a fraction

OpenStudy (anonymous):

it was a fraction

OpenStudy (unklerhaukus):

\[5^1=5\]

OpenStudy (lgbasallote):

hint \[\frac{a^4}{a^1} = a^{4-1}\]

OpenStudy (anonymous):

1/ (5^4)(5) = 1/ 3125

OpenStudy (anonymous):

bring down 5^-4 to make it positive

jhonyy9 (jhonyy9):

this is equal 5^(-5)

jhonyy9 (jhonyy9):

or 1/5^5

OpenStudy (anonymous):

thx

jhonyy9 (jhonyy9):

yw good luck bye

jhonyy9 (jhonyy9):

but i hope so much that you know the property of terms with exponents

jhonyy9 (jhonyy9):

like a^b * a^c =a^(b+c)

jhonyy9 (jhonyy9):

or a^(-1)=1/a